OpenStudy (happykiddo):

Two violinists are trying to tune their instruments in an orchestra. One is producing the desired frequency of 440.0 Hz. The other is producing a frequency of 453 Hz. By what percentage should the out-of-tune musician change the tension in his string to bring his instrument into tune at 440.0 Hz?

OpenStudy (jtug6):

So the first thing you need to ask is what is the difference of 440hz from 453hz?

OpenStudy (happykiddo):

13

OpenStudy (jtug6):

Right. So now the question is what percentage so we can write out an equation to look like this: 400 * x = 13 where x is the percentage needed to get us 13 hz

OpenStudy (jtug6):

Oops my bad i just woke up i mean 453*x=13

OpenStudy (jtug6):

So what would x equal then?

OpenStudy (happykiddo):

so are you saying to divide 13 by 400, and then multiply by 100 to get the percent

OpenStudy (jtug6):

Yes, but 453 not 400 my mistake :P

OpenStudy (jtug6):

And so now you know it's around 2.9 percent so reverse check it: 453*0.029 = around 13 which 453-13 = 440 which is what we were looking for

OpenStudy (happykiddo):

Alright thanks I get it! I tried using some explanation on yahoo and it was crap. Have a nice day.

OpenStudy (jtug6):

Mhm. Just be careful w/the answer cause it may want it rounded to a whole number/tenth ect :P

OpenStudy (happykiddo):

okay

OpenStudy (happykiddo):

2.9 was wrong the answer is 5.7%

OpenStudy (jtug6):

I'm sorry. That should be the way though. Even using a percentage difference calculator is gives 2.9 percent so i really dont know how else that could be wrong

OpenStudy (jtug6):

Does it give an explanation?

OpenStudy (happykiddo):

It gives no explanation but there is that yahoo answers link that I couldn't understand. Maybe you could take a look at it. https://answers.yahoo.com/question/index?qid=20100502081857AAQQ2AG

OpenStudy (jtug6):

Ok well I'm terribly sorry but ive never seen an equation like that. However, it should be (440/453)^2 * 100 then 100 - the answer you get from the left which equals the 5.7%.

OpenStudy (happykiddo):

alright thanks!